A 1.86 kg book in space has a weight of 5.08 N.
What is the value of gravitational field at
that location?
1
Expert's answer
2017-11-30T16:03:07-0500
Newton's second law says, that F =mg
Then g= F/m= (5.08 N)/(1.86 kg)=2.59 m/s^2
Standard gravity (g) = 2.59 meter per second square.